OpenStudy (anonymous):

Evaluate: [(40 + 5) − 32]^2^ ÷ 9 • 2

OpenStudy (anonymous):

Combine whats inside first...45-32 = 13 13^2 = 169 / 18 = 9.388888889

OpenStudy (anonymous):

9.39